lauragabrielle:

A 10" diameter pumpkin pie is cut into six equal servings. What is the approximate area of the top of each piece of pie? 3.33π inches 4.17π inches 16.66π inches

Vocaloid:

area = pi*r^2 area of one slice = (pi*r^2)/6 since there are 6 equal slices to find radius, divide the diameter 10 by 2, then plug that into the area of one slice equation since the solution are in terms of pi, simplify everything in the equation except pi, leave your answer in terms of pi